Riverbed Corporation reported net income of $259,140 in 2020 and had 52,100 shares of common stock outstanding throughout the year. Also outstanding all year were 4,800 shares of cumulative preferred stock, each convertible into 2 shares of common. The preferred stock pays an annual dividend of $5 per share. Riverbed’s tax rate is 40%.

Compute Riverbed’s 2020 diluted earnings per share. (Round answer to 2 decimal places, e.g. 3.55.)

Diluted earnings per share $
Answer #1

Basic earning per share = (259140-4800*5)/52100 = 4.51 per share

Diluted earning per share = Adjusted net income/Adjusted shares = 259140/(4800*2+52100) = 4.20 per share